OpenStudy (anonymous):

Mike went to a football game with his family. They bought 5 boxes of popcorn, 3 programs and spent $41. Sean’s family went to the same game, bought 3 boxes of popcorn, 2 programs and spent $26. How much does each program cost at the game? $6 $4 $7 $5

OpenStudy (anonymous):

Let there are x boxes of popcorn, and there be y programs: So, 5x + 3y = 41 3x + 2y = 26 Multiply first equation by 3 and second equation by 5: 15x + 9y = 123 15x + 10y = 130 Just subtract them you will have the y which will be your answer.. Can you do it??

OpenStudy (anonymous):

so you subtract the x from the 123 and 130?

OpenStudy (anonymous):

I think x and x will get can cancel.. you will have only on left hand side..

OpenStudy (anonymous):

y...

OpenStudy (anonymous):

is the first one y=12?

OpenStudy (anonymous):

What do you get when you subtract 123 from 130??

OpenStudy (anonymous):

ohhh you subtract those two numbers okay cool its 7